OpenStudy (anonymous):

What is the solution to the following system of equations? 2x + 3y = 6 x + y = 4

OpenStudy (aayushi.somani):

find the value of x by ur 2nd equation..! x=4-y and put it in the 1st equation-- 2(4-y)+3y=6 8-2y+3y=6 y=-2 now use the value of y in the equation for x-- x=4-y x=2+2 x=4